OpenStudy (anonymous):

The lengths (in feet) of two sides of a regular quadrilateral are represented by the expressions 8x-6 and 4x+22. Find the length of a side of the quadrilateral help me with the work too because I have no clue what I'm doing. Thank you (-:

OpenStudy (anonymous):

hellllo

OpenStudy (anonymous):

@brieculver

OpenStudy (anonymous):

are u ready or no?

OpenStudy (anonymous):

A regular quadrilateral is a sqaure because the angles are equal and the sides are equal. so 8x - 6 = 4x + 22 8x = 4x + 22 + 6 8x = 4x + 28 4x = 28 x = 7

OpenStudy (anonymous):

thank you @SADIQ6205